I have to solve for B and I have no clue how to do this
The teacher said it can be done in 1-2 steps
AB-CD = BD+AC

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(48535)

AB-CD = BD+AC
--------------------
Get the "B-Terms" together:
AB - BD = AC + CD
------------------------
Factor out the "B":
B(A-D) = AC + CD
----------------------
Solve for B:
B = (AC+CD)/(A-D)
=========================
Cheers,
Stan H.

Answer by Alan3354(21583)

I have to solve for B and I have no clue how to do this
The teacher said it can be done in 1-2 steps
AB-CD = BD+AC
--------------------
You can only solve for B in terms of the other 3 variables.
Add CD to both sides
AB = BD+AC+CD
Get all terms with B on the same side by subtracting BD from both sides
AB-BD = AC+CD
Factor the terms
B*(A-D) = C*(A+D)
Divide by (A-D) to get B by itself
--------------------
B = C*(A+D)/(A-D)
That's as far as you can go.